OpenStudy (jjenk):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). Help Please

jimthompson5910 (jim_thompson5910):

f(x) = -4x + 7 f(g(x)) = -4( g(x) ) + 7 ... replace every x with g(x) f(g(x)) = -4( 10x-6 ) + 7 ... replace every g(x) on the right side with 10x-6 f(g(x)) = -4*10x-4*(-6 ) + 7 f(g(x)) = -40x + 24 + 7 f(g(x)) = -40x + 31